OpenStudy (anonymous):

In mixing a bug spray, a 40% solution of the chemical is mixed with an 85% solution to get 20 litres of a 60% solution. How much of each solution is required?

OpenStudy (kropot72):

The amount of chemical in the 20 litres of the 60% solution is 20 * 0.6 = 12 litres. Let A be the amount of 40% solution needed and let B be the amount of 85% solution needed. We can now write the following equation based on the amounts of chemical: \[\large 12=0.4A+0.85B\ ............(1)\] We know that the combined amounts of 40% and 85% solutions needed to make the final 60% solution must equal 20 litres. Therefore we can write: \[\large A+B=20\ ........(2)\] and rearranging (2) we get: \[\large B=20-A\ .............(3)\] Now we can plug the expression for B given in equation (3) into equation (1) to give: \[\large 12=0.4A+0.85(20-A)\ .........(4)\] Equation (4) can be simplified to give: \[\large 0.45A=5\ ............(5)\] Can you now solve equation (5) to find the value of A, and then substitute that value in equation (2) to find the value of B?

OpenStudy (anonymous):

I must be really stupid, my mind isn't working and it doesn't make sense 1 can solve 5

OpenStudy (anonymous):

can't

OpenStudy (kropot72):

\[\large A=\frac{5}{0.45}=you\ can\ calculate\]

OpenStudy (anonymous):

100/9 or 11.111111

OpenStudy (anonymous):

b=20-100/9= 80/9 or 8.888888889

OpenStudy (kropot72):

Yes, you are correct. the exact answers are: \[\large A=11\frac{1}{9}\] \[\large B=8\frac{8}{9}\]